### v2

# Quantum emitters in silicon: Purcell-enhanced lifetime modulation as a probe of local refractive index changes

Abstract
Quantum emitters embedded in silicon photonic cavities offer a powerful and scalable platform for label-free refractive index sensing at the nanoscale. We propose and theoretically analyze a novel sensing mechanism based on Purcell-enhanced modulation of the emitter’s spontaneous emission lifetime, enabling detection of refractive index changes via time-correlated single-photon counting (TCSPC). In contrast to traditional resonance-shift sensors, our approach exploits the lifetime sensitivity to variations in the local density of optical states (LDOS), providing an intensity-independent, spectrally unresolvable, and CMOS-compatible sensing modality. We derive analytical expressions linking refractive index perturbations to relative lifetime shifts and identify an optimal off-resonance operation regime, where the lifetime response becomes linear and maximally sensitive to small perturbations. Parametric evaluation of the analytical expressions shows that, for moderate-quality photonic cavities ($Q$ = 105–107), this method enables refractive index detection limits as low as 10-9 RIU — competitive with or even outperforming state-of-the-art plasmonic and microresonator sensors, yet requiring significantly simpler instrumentation. Furthermore, we show that long-lived emitters such as T-centers in silicon provide a unique advantage, allowing sub-nanosecond lifetime shifts to be resolved with standard TCSPC systems. Although room-temperature operation of quantum emitters in silicon has yet to be demonstrated, our results lay the theoretical foundation for scalable, room-temperature, quantum-enabled sensing of refractive index changes — eliminating the need for spectral resolution and cryogenic infrastructure.
I Introduction
[…]
However, their practical use at room temperature remains constrained by phonon-induced dephasing and non-radiative recombination, which broaden and suppress ZPL emission. To address these limitations, we propose a fundamentally different transduction approach: lifetime-based sensing via modulation of the local density of optical states (LDOS). Unlike conventional schemes based on spectral shifts, this method relies on variations in the emitter’s spontaneous emission rate induced by its electromagnetic environment. Such a detection scheme is inherently more robust to thermal broadening and inhomogeneous line shifts, and remains functional even when the ZPL becomes spectrally unresolved. In other words, even if the emitter rapidly loses coherence (i.e., short $T_{2}$) and exhibits a broad spectral line, the excited-state population may still decay on a much longer timescale ($T_{1}$), yielding a measurable lifetime — the quantity we exploit for sensing. This approach extends beyond T-centers to a broader class of T-like color centers, including group III — carbon complexes such as B–C, Al–C, and In–C [7, 8]. These centers are predicted to exhibit similar spin and optical properties, with higher structural symmetry (e.g., $C_{2v}$), improved radiative efficiencies, and greater thermal stability — making them promising candidates for scalable integration with silicon photonics.
The viability of this lifetime-based sensing scheme is further supported by decades of work on quantum spin systems in silicon. For instance, phosphorus donor nuclear spins in isotopically purified silicon have shown coherence times exceeding 39 minutes at room temperature [2], while 29Si nuclear spin qubits maintain coherence even without isotopic enrichment [3], affirming silicon’s status as an ultra-low-noise host for quantum systems. Analogous temperature studies of color centers such as SiV in diamond provide insights into emission linewidths: homogeneous broadening increases with temperature following a $T^{3}$ dependence, leading to ZPL broadening from $\sim$0.1–0.6 nm at cryogenic temperatures to $\sim$1.3–2 nm at room temperature [9]. A similar trend is anticipated for silicon-based centers, though experimental verification remains pending. Although room-temperature operation of quantum emitters in silicon remains unproven, no conclusive physical limitation has been established, suggesting potential for future advancement through photonic or material engineering.
In light of these considerations, our proposed sensing scheme — based on photonic cavity-enhanced lifetime modulation of optically active color centers [10, 11, 12, 13, 14] — remains viable under room-temperature conditions. By decoupling sensing performance from spectral resolution constraints, it offers a robust and scalable platform for refractive index detection and single-molecule interaction studies. It is inherently compatible with the future discovery and optimization of silicon-based quantum emitters.
Moreover, this approach occupies a distinctive position among existing quantum sensing methodologies. Unlike nitrogen-vacancy (NV) centers in diamond — which require intricate spin-state manipulation and suffer from integration challenges — our method is fully compatible with established silicon photonic fabrication technologies. While fluorescence lifetime modulation via surrounding refractive index modification has previously been explored using NV centers in diamond [15], those approaches typically do not explicitly incorporate the photonic environment via a cavity with defined quality factor $Q$. In contrast, our model analytically links refractive index changes $\Delta n$ to spontaneous emission lifetime through the Purcell factor, which depends on both $\Delta n$ and $Q$. This introduces a new degree of tunability and physical insight, positioning this method as a fundamentally distinct sensing strategy. In addition, while NV-based quantum sensors have demonstrated lifetime-based sensing in magnetometry and microfluidic environments [16, 17, 18], the rigid crystalline nature of diamond hinders their monolithic integration with scalable photonic circuitry. By leveraging T-center and T-like quantum emitters embedded in silicon waveguides, our platform enables site-controlled integration with photonic cavities on a CMOS-compatible substrate. Crucially, it allows the detection of LDOS-modulated lifetime shifts without the need for coherent spin control. This makes the scheme highly attractive for practical, room-temperature sensing and single-molecule interaction detection in compact, scalable devices — opening a viable path toward quantum-enabled photonic sensing beyond the limitations of spectral resolution and cryogenic operation.
II Theoretical framework and results
The spontaneous emission lifetime of a quantum emitter coupled to an optical cavity is governed by the Purcell factor ($F_{P}$[19]:
$F_{P}(\vec{r},\theta)=\frac{3}{4\pi^{2}}\left(\frac{\lambda}{n}\right)^{3}\frac{Q}{V}\left|\frac{\vec{E}(\vec{r})\cdot\vec{d}}{|\vec{E}_{\text{max}}||\vec{d}|}\right|^{2},$ (1)
where $\lambda$ is the emission wavelength, $n$ is the refractive index of the cavity medium, $Q$ is the quality factor of the cavity, and $V$ is the effective mode volume. The term $\vec{E}(\vec{r})$ is the local electric field of the cavity mode at the emitter position $\vec{r}$, $\vec{E}_{\text{max}}$ is the maximum field amplitude in the cavity, and $\vec{d}$ is the unit vector along the emitter’s dipole moment. The angle $\theta$ denotes the angle between the dipole moment $\vec{d}$ and the local electric field vector $\vec{E}(\vec{r})$, such that $\left|\vec{E}(\vec{r})\cdot\vec{d}\right|^{2}=|\vec{E}(\vec{r})|^{2}\cos^{2}\theta$. This expression accounts for both spatial and orientational mismatch between the emitter and the cavity field, reducing the effective Purcell factor in cases where the emitter is not ideally positioned or aligned.
We start with the ideal case in which the emitter dipole is perfectly aligned with the cavity field, and the emitter is located at the field maximum. In this case, $\theta=0$ and $\vec{E}(\vec{r})=\vec{E}_{\text{max}}$, so Eq. (1) simplifies to:
$F_{P}=\frac{3}{4\pi^{2}}\left(\frac{\lambda}{n}\right)^{3}\frac{Q}{V}.$ (2)
The modified spontaneous emission lifetime $\tau$ is then defined as:
$\tau=\frac{\tau_{0}}{F_{P}},$ (3)
where $\tau_{0}$ is the intrinsic spontaneous emission lifetime of the emitter in free space or in an unstructured (homogeneous) dielectric. We assume that the optical cavity used to modify the spontaneous emission lifetime of the emitter is optimized to achieve the maximum LDOS and Purcell enhancement. Local changes in the refractive index induce a spectral detuning of the cavity resonance. This detuning reduces the LDOS at the emitter’s location, thereby suppressing the spontaneous emission rate and resulting in an increase in the emitter’s lifetime. This lifetime shift can be detected using time-correlated single-photon counting (TCSPC), forming the basis for a label-free sensing modality. Given this, we estimate the effect of a refractive index change on the spontaneous emission lifetime of a quantum emitter coupled to a photonic cavity. A local change in the refractive index alters the effective index of the cavity mode, $n_{\mathrm{eff}}$, leading to a shift in cavity’s resonance angular frequency $\omega_{\mathrm{cav}}$. In the regime of small perturbations, the shift in the resonance frequency can be expressed using first-order perturbation theory as:
[…]
where $\Delta\varepsilon(\vec{r})=2n(\vec{r})\Delta n(\vec{r})$ is the permittivity perturbation induced by the local refractive index change $\Delta n(\vec{r})$, and $\vec{E}(\vec{r})$ is the cavity mode electric field. By assuming $\Delta\varepsilon=2n\Delta n\approx 2n_{\mathrm{eff}}\Delta n$, we can then simplify Eq. (4) to:
$\Delta\omega=\omega_{\mathrm{cav}}\frac{\Delta n}{n_{\mathrm{eff}}}.$ (5)
Depending on cavity type, this detuning lowers the spectral overlap between the emitter (e.g., T-center) and the cavity mode, reducing the Purcell factor. The Purcell factor $F_{P}$ under detuning follows a Lorentzian dependence:
$F_{P}(\Delta\omega)=\frac{F_{P,\mathrm{max}}}{1+\left(\frac{2Q\Delta\omega}{\omega_{\mathrm{cav}}}\right)^{2}},$ (6)
where $\Delta\omega=\omega_{\mathrm{emitter}}-\omega_{\mathrm{cav}}$ is the detuning. If the cavity is on-resonance in the initial state: $\Delta\omega=0$, and $F_{P}=F_{P,\mathrm{max}}$, and lifetime is at its minimum ($\tau=\tau_{0}/F_{P}$). Then, a small change in $F_{P}$ leads to the following relative lifetime change of the emitter:
$\frac{\Delta\tau}{\tau}\approx\left(\frac{2Q\Delta n}{n_{\mathrm{eff}}}\right)^{2}.$ (7)
For parametric evaluation of Eq. (7), we assume the following values: $Q=10^{5}$, $10^{6}$, or $10^{7}$ and $n_{\mathrm{eff}}=2.5$. Simulated relative lifetime change $\Delta\tau/\tau$ versus refractive index change $\Delta n$ comparing the TCSPC approach with the state-of-the-art plasmonic detection limit is shown in Figure 1, where Figure 1a features linear scale for the $\Delta\tau/\tau$ ranging from 0 to 10% and logarithmic scale for the $\Delta n$, while Figure 1b — both in logarithmic scale.
Figure 1: Simulated lifetime change ($\Delta\tau/\tau$) versus refractive index change ($\Delta n$): (a) linear scale for $\Delta\tau/\tau$ from 0 to 10% and logarithmic scale for $\Delta n$; (b) logarithmic scale for $\Delta\tau/\tau$ from $10^{-1}$ to $10^{5}$ and logarithmic scale for $\Delta n$.
According to Figure 1, in cases of the cavity with $Q=10^{5}$ in the on-resonance state, achieving the state-of-the-art sensitivity of the plasmonic measurements ($10^{-6}$ RIU) requires the ability of measuring the relative lifetime changes of less than 1%. Solving Eq. (7) for $\Delta n$ and substituting these values shows that we are potentially sensitive to refractive index changes as small as $\Delta n\approx 4.68\times 10^{-7}$ ($Q=10^{5}$), which may compete with or even surpass the sensitivity of state-of-the-art plasmonic devices. In general, for a photonic cavity with $Q=10^{5}$, achieving a relative lifetime change $\Delta\tau/\tau$ below 0.64% (intersection of the yellow solid curve and red dashed line in Figure 1) would enable the proposed system to surpass the $10^{-6}$ RIU sensitivity limit typical of plasmonic sensors.
However, at resonance ($\Delta\omega=0$), the Purcell factor is at its maximum, but its derivative with respect to detuning is zero. Therefore, in the immediate vicinity of $\Delta n=0$ (i.e., near the point $n_{\mathrm{eff}}$, see Figures 1a and 2), the dependence of $\Delta\tau/\tau$ on $\Delta n$ becomes non-linear (specifically, quadratic; see Eq. (7)), making it unsuitable for applications that require quantitative detection of small refractive index changes.
Figure 2: Simulated normalized Purcell factor versus refractive index change $\Delta n$: (a) linear scale for $\Delta n$ from $-10^{-4}$ to $+10^{-4}$; (b) log scale for $\Delta n$ from $10^{-9}$ to $10^{-3}$.
A more effective approach is to shift the emitter-cavity system off-resonance, to the point where the slope of the Lorentzian Purcell factor is maximized — i.e., at detuning by $\Delta\omega=\pm\kappa/2=\pm\omega_{\mathrm{cav}}/(2Q)$ where $\kappa$ is the cavity decay rate. This $\Delta\omega$ detuning relates to the following required shift in the refractive index from the on-resonance point, $n_{\mathrm{shift}}=n_{0}-n_{\mathrm{eff}}=\pm n_{\mathrm{eff}}/2Q$, where $n_{0}$ is the off-resonance point we want to work around: $n_{0}=n_{\mathrm{eff}}\pm n_{\mathrm{eff}}/2Q=n_{\mathrm{eff}}(1\pm 1/2Q)$. Here, the relative lifetime change becomes linearly dependent on $\Delta n$ for small perturbations, enabling maximum sensitivity and restoring a wide, usable dynamic range (see Figure 2). At these points (marked by vertical black lines in Figure 2a for $Q=10^{5}$ and by color dots in Figure 2b for all curves), the lifetime of quantum emitters is not at its minimum, but its sensitivity to detuning is maximal. This trade-off introduces a moderate ($\sim$50%) reduction in Purcell enhancement but enables a linear, tunable sensing regime suitable for detecting small $\Delta n$ changes. This strategy also introduces a modest reduction in intensity (since emission rate is lower), but dramatically improves the usable dynamic range. The reduction in intensity can be mitigated by better photodetectors, longer integration, and higher emitter density. Especially for long-lived emitters like T-centers, this intensity loss is well tolerated, and the resulting linear response regime is highly advantageous for sensing applications.
We next derive the expression for the detection limit related to these optimal detection points for each $Q$ value, $\Delta n_{\mathrm{opt}}$. The emission lifetime $\tau$ at the shifted point $n_{0}$ can be written as:
$\tau(n_{0})=\frac{\tau_{\mathrm{eff}}}{1+\left(2Q\frac{(n_{0}-n_{\mathrm{eff}})}{n_{\mathrm{eff}}}\right)^{2}},$ (8)
where $\tau_{\mathrm{eff}}$ is the lifetime of the quantum emitter in the on-resonance state. In the approximation of $\Delta n\ll n_{\mathrm{shift}}$, the emission lifetime change around $n_{0}$ can be expressed as follows:
$\frac{d\tau(n_{0})}{dn_{0}}$ $=-\frac{\tau_{\mathrm{eff}}}{\left[1+\left(2Q\frac{(n_{0}-n_{\mathrm{eff}})}{n_{\mathrm{eff}}}\right)^{2}\right]^{2}}\cdot 2\cdot 2Q\frac{(n_{0}-n_{\mathrm{eff}})}{n_{\mathrm{eff}}}\cdot\frac{2Q}{n_{\mathrm{eff}}}$
$=-4\tau_{0}\frac{Q^{2}}{n_{\mathrm{eff}}^{2}}n_{\mathrm{shift}}$
=±$=\pm 2\tau_{0}\frac{Q}{n_{\mathrm{eff}}}=\pm\tau_{\mathrm{eff}}\frac{Q}{n_{\mathrm{eff}}},$ (9)
here $\tau_{0}=\tau(n_{0})=0.5\tau(n_{\mathrm{eff}})=0.5\tau_{\mathrm{eff}}$. In the vicinity of $n_{0}$, we can write:
$\tau(n_{0}+\Delta n)$ $=\tau(n_{0})+\frac{d\tau(n_{0})}{dn_{0}}\Delta n$
$=\tau_{0}\pm 2\tau_{0}\frac{Q}{n_{\mathrm{eff}}}\Delta n$
$=\tau_{0}\left(1\pm 2\frac{Q}{n_{\mathrm{eff}}}\Delta n\right).$ (10)
From Eq. (10) it follows that:
$\Delta\tau=\pm 2\tau_{0}\frac{Q}{n_{\mathrm{eff}}}\Delta n,$ (11)
and consequently:
$\left|\frac{\Delta\tau/\tau_{0}}{\Delta n}\right|=2\cdot\frac{Q}{n_{\mathrm{eff}}}.$ (12)
Eq. (12) gives the slope of the Lorentzian response curve at the point of maximum sensitivity $n_{0}$. According to Eq. (12), for a typical effective refractive index of $n_{\mathrm{eff}}=2.5$, the sensitivity slopes are depicted in Table 1.
Table 1: Sensitivity slopes as a function of the $Q$ factor of the cavity.
Q $|\Delta\tau/\tau|$ per $\Delta n$ (RIU-1)
$10^{5}$ $8\times 10^{4}$
$10^{6}$ $8\times 10^{5}$
$10^{7}$ $8\times 10^{6}$
These values represent the slope of the relative lifetime change per unit refractive index change, measured at the point of optimal detuning ($\Delta\omega=\kappa/2$). This slope enables conversion between $\Delta n$ and the required measurable $\Delta\tau/\tau_{0}$ for a given $Q$. Eq. (12) can be re-written for $\Delta n_{\text{opt}}$ as:
$\Delta n_{\text{opt}}=\frac{\left(\Delta\tau/\tau_{0}\right)}{2Q}\cdot n_{\text{eff}}.$ (13)
Eq. (13) estimates the minimum detectable $\Delta n_{\text{opt}}$ for a given temporal resolution. According to Eq. (13), the detection limit of the proposed refractive index sensing approach in the off-resonance regime, $\Delta n_{\text{opt}}$, is limited by the $Q$ value of the photonic cavity and the resolution capability $\Delta\tau/\tau_{0}$ of the TCSPC system used. For instance, if a TCSPC system can resolve a 3% change in lifetime ($\Delta\tau/\tau_{0}=0.03$), then the minimum detectable $\Delta n_{\text{opt}}$ would be approximately $3.75\times 10^{-7}$ RIU for $Q=10^{5}$, $3.75\times 10^{-8}$ RIU for $Q=10^{6}$, and $3.75\times 10^{-9}$ RIU for $Q=10^{7}$, assuming $n_{\text{eff}}=2.5$. In addition, we calculated the linear dynamic range (LDR) for each $Q$ value as the span of refractive index changes $\Delta n$ over which the relative lifetime change $\Delta\tau/\tau_{0}$ remains approximately linear (with a deviation less than $\pm$10from linear behavior) around the $n_{0}$ point. These data are summarized in Table 2.
Table 2: Minimum detectable $\Delta n_{\text{opt}}$ and LDR width as a function of $Q$.
Q Minimum $\Delta n_{\text{opt}}$ [RIU] LDR width [RIU]
$10^{5}$ $3.75\times 10^{-7}$ $1.11\times 10^{-5}$
$10^{6}$ $3.75\times 10^{-8}$ $1.11\times 10^{-6}$
$10^{7}$ $3.75\times 10^{-9}$ $1.11\times 10^{-7}$
At this point we must address the consequences of the idealizations made when transitioning from Eq. (1) to Eq. (2). In practice, achieving perfect spatial alignment and dipole orientation of quantum emitters in silicon photonic cavities is highly challenging with current technology. For example, standard techniques such as ion implantation followed by thermal annealing [20] result in emitters that are distributed in space with a resolution no better than 20 nm and possess stochastic dipole orientations relative to the cavity mode field. At the same time, the Purcell factor $F_{P}$ is strongly dependent not only on the emitter’s spectral resonance with the cavity mode but also on its position within the cavity and the relative orientation of its dipole moment with respect to the local electric field [21]. As a result, Eq. (12), which describes the slope of the Lorentzian lifetime response curve at the point of maximum sensitivity $n_{0}$, reflects an idealized scenario assuming maximum Purcell coupling — i.e., the emitter is located at the field antinode and its dipole is perfectly aligned with the local field vector. In realistic implementations, however, spatial and orientational averaging must be taken into account.
To capture this deviation, we introduce an effective coupling efficiency factor $\eta_{\mathrm{eff}}\in[0,1]$, which quantifies the reduction of the average Purcell factor due to positional and angular mismatch. This factor can be expressed as:
$\eta_{\mathrm{eff}}=\left\langle\cos^{2}\theta\right\rangle\cdot\frac{\left\langle\left|\vec{E}(\vec{r})\right|^{2}\right\rangle}{\left|\vec{E}_{\mathrm{max}}\right|^{2}},$ (14)
where $\left\langle\cos^{2}\theta\right\rangle=\int_{0}^{\pi}\cos^{2}\theta\sin\theta\,d\theta\,/\,\int_{0}^{\pi}\sin\theta\,d\theta=\frac{1}{3}$ represents the average projection of the dipole on the field direction for isotropic orientation in three dimensions, and $\left\langle|\vec{E}(\vec{r})|^{2}\right\rangle/|\vec{E}_{\mathrm{max}}|^{2}$ accounts for the average field intensity experienced by the emitters relative to the maximum. If the cavity mode field has the form $E(x)\sim\cos(kx)$, then the intensity is given by $|E(x)|^{2}\sim\cos^{2}(kx)=\left(1+\cos(2kx)\right)/2$. When averaging over the spatial coordinate $x$, we obtain $\langle\cos^{2}(kx)\rangle=1/2$. According to Eq. (14), this yields $\eta_{\mathrm{eff}}\sim 1/6$ as a reasonable estimate. Taking this factor into account, the realistic slope of the Lorentzian lifetime response curve becomes:
$\left|\frac{\Delta\tau/\tau_{0}}{\Delta n}\right|_{\mathrm{realistic}}=\eta_{\mathrm{eff}}\cdot\left(\frac{2Q}{n_{\mathrm{eff}}}\right)\approx\frac{1}{6}\cdot\left(\frac{2Q}{n_{\mathrm{eff}}}\right),$ (15)
which can be interpreted as the effective sensitivity of the system in realistic experimental conditions. While this reduction is not negligible, it does not negate the viability of the sensing scheme. On the contrary, it highlights the importance of modeling the average response and sets a benchmark for achievable performance using ensembles of implanted quantum emitters. Furthermore, the inherent sensitivity of long-lived emitters (such as T-centers) combined with the robustness of TCSPC techniques ensures that even a fraction of the ideal Purcell response can yield detectable and quantifiable lifetime shifts for practical sensing applications. Notably, this reduction in effective sensitivity due to spatial and orientational averaging can be partially compensated by employing cavities with one order higher $Q$, thereby restoring comparable performance.
It is worth noting that while Eq. (14) provides a useful approximation for ensemble coupling efficiency under isotropic and homogeneous assumptions, an alternative form involving an explicit sum over emitter positions and orientations can be written as:
$\eta_{\text{eff}}=\frac{1}{N}\sum_{i=1}^{N}\cos^{2}\theta_{i}\cdot\frac{|\vec{E}(\vec{r}_{i})|^{2}}{|\vec{E}_{\text{max}}|^{2}},$ (16)
where $\theta_{i}$ is the angle between the $i$-th emitter’s dipole moment and the local electric field, and $\vec{r}_{i}$ is its spatial position. This formulation reduces to Eq. (14) when spatial and angular averaging is applied under the assumption of uniform emitter distribution and isotropic orientation. Such an emitter-resolved form can be particularly useful in the context of precise emitter engineering — for example, when emitters are deterministically placed at specific field maxima or aligned to maximize coupling. In the limit of a continuous distribution $\rho(\vec{r})$, the average coupling efficiency may be expressed as:
[…]
From these considerations, it is clear that the proposed sensing mechanism can be considered as future-forward, targeting best-in-class performance, as it can potentially resolve $\Delta n$ changes down to $10^{-9}$ RIU. Even lower values can be readily achievable for cavities with $Q$-factors $>10^{7}$, achievable via state-of-the-art silicon photonics cavity designs [22]. This corresponds to a minimum detectable refractive index shift of $<10^{-9}$ RIU, outperforming plasmonic counterparts and unlocking single-molecule sensitivity in compact, integrated devices.
Another practical advantage of operating the sensor off-resonance is the significant relaxation of the technical requirements on the time-resolved detection system. At exact resonance ($\Delta\omega=0$), the relative lifetime change in response to small refractive index shifts can be extremely low (see Figure 1a) — often below 1% — which necessitates ultra-high-resolution TCSPC instrumentation. In contrast, at the optimal detuning point ($\Delta\omega=\kappa/2$), the slope of the Purcell factor with respect to detuning is maximized, enabling much larger and more easily measurable lifetime changes for the same $\Delta n$. As a result, the proposed off-resonance sensing strategy enables high sensitivity while maintaining a practical detection limit for the TCSPC system.
While higher-$Q$ cavities generally enhance sensitivity, they do not always translate into stricter requirements for the TCSPC system—especially when operating off-resonance. At higher $Q$ values, the Purcell effect becomes stronger, leading to shorter emitter lifetimes in the initial state. However, if the system is intentionally operated at the point of maximum slope of the Purcell curve (i.e., off-resonance at $\Delta\omega=\kappa/2$), the relative lifetime changes become significantly larger and more linear, even for small refractive index shifts. This approach reduces the need for ultra-high temporal resolution, as detecting a few percent change in lifetime (instead of sub-percent) is well within the capability of modern TCSPC systems. At the same time, increasing $Q$ introduces other challenges, such as tighter fabrication tolerances, greater thermal sensitivity, and more demanding integration procedures. The choice of $Q$ should therefore balance sensitivity, fabrication feasibility, and instrumentation requirements.
III Advantage of long-lived emitters like T-centers
One of the key technical challenges in high-$Q$ cavity-enhanced sensing is the requirement for very high time resolution in TCSPC systems, especially when using fast-emitting fluorophores like organic dyes ($\tau\sim 3.5$ ns). In those cases, small relative lifetime changes (e.g., 0.1–0.3%) correspond to absolute shifts in the range of 3–10 ps — demanding the use of specialized ultrafast detection equipment. However, T-centers in silicon offer a unique advantage due to their long natural lifetimes, often in the microsecond range. For instance, a T-center with a cavity-unaffected (uncoupled) lifetime of 1 $\mu$[14], when coupled to a photonic cavity with a moderate Purcell factor (e.g., $F_{P}=5$), will exhibit a reduced lifetime of about 200 ns. In this regime, even a 0.1% lifetime change corresponds to a shift of 200 ps — easily detectable with standard TCSPC systems. This greatly relaxes the instrumentation requirements and enables detection of ultra-small refractive index changes, making T-centers a highly attractive emitter platform for integrated biosensing using cavity-enhanced lifetime detection. Figure 3 shows a comparison of effective lifetimes $\tau_{\mathrm{eff}}$ of a typical fluorophore (3.5 ns intrinsic) and a T-center with an uncoupled lifetime of 1 $\mu$s, as a function of Purcell factor. T-centers maintain a large $\tau_{\mathrm{eff}}$ ($\sim$1 ns) even under high Purcell enhancement ($F_{P}=10^{3}$), easing TCSPC requirements.
Figure 3: Comparison of effective lifetimes $\tau_{\mathrm{eff}}$ of a typical fluorophore ($\sim$3.5 ns intrinsic) and a T-center with an uncoupled lifetime of 1 $\mu$s, as a function of Purcell factor.
In the context of off-resonance sensing, where the system is operated at the point of maximum slope in the Purcell response curve, a relative change in lifetime of a few percent (e.g., 3%) is both realistic and desirable. This linear regime simplifies signal interpretation and reduces the burden on instrumentation. The absolute lifetime change required for detection is approximately $\Delta\tau=0.03\cdot\tau_{\mathrm{eff}}$, where $\tau_{\mathrm{eff}}$ is the cavity-enhanced lifetime. This requirement depends on the emitter type and the cavity-enhanced lifetime $\tau_{\mathrm{eff}}$, which is reduced from the intrinsic value via the Purcell effect as shown in Table 3.
Table 3: Estimated lifetime and $\Delta\tau$ requirements for fluorescent dyes (3.5 ns) and T-centers (1 $\mu$s) under various $Q$ values assuming off-resonance sensing.
Q Fluorescent Dyes T-Centers
$10^{5}$ Est. $F_{P}$: $\sim$10 $\tau_{\mathrm{eff}}$: $\sim$0.35 ns $\Delta\tau$ (3%): $\sim$10.5 ps Required TCSPC: $\sim$10–20 ps (high-end) Est. $F_{P}$: $\sim$10 $\tau_{\mathrm{eff}}$: $\sim$100 ns $\Delta\tau$ (3%): $\sim$3 ns Required TCSPC: $\sim$2–5 ns (standard)
$10^{6}$ Est. $F_{P}$: $\sim$50 $\tau_{\mathrm{eff}}$: $\sim$0.07 ns $\Delta\tau$ (3%): $\sim$2.1 ps Required TCSPC: $<$10 ps (cutting-edge) Est. $F_{P}$: $\sim$50 $\tau_{\mathrm{eff}}$: $\sim$20 ns $\Delta\tau$ (3%): $\sim$600 ps Required TCSPC: $\sim$500 ps–1 ns (standard)
$10^{7}$ Est. $F_{P}$: $\sim$100 $\tau_{\mathrm{eff}}$: $\sim$0.035 ns $\Delta\tau$ (3%): $\sim$1.0 ps Required TCSPC: $<$5 ps (advanced only) Est. $F_{P}$: $\sim$100 $\tau_{\mathrm{eff}}$: $\sim$10 ns $\Delta\tau$ (3%): $\sim$300 ps Required TCSPC: $\sim$200–500 ps (standard)
As shown in Table 3, fluorescent dyes require very high timing resolution (20 ps or better) for high-$Q$ cavities, making them challenging to use without ultrafast detection systems. In contrast, T-centers, even when strongly coupled to high-$Q$ cavities, maintain effective lifetimes in the tens to hundreds of nanoseconds. This translates into much larger absolute lifetime changes, which can be resolved using standard TCSPC systems with time resolution in the hundreds of picoseconds to a few nanoseconds. This distinction makes T-centers a highly attractive choice for scalable and practical high-sensitivity sensing.
IV Comparison with other approaches based on resonance wavelength shifts
In addition to comparing our approach with plasmonic measurements, it is also important to evaluate it against other sensing methods based on resonance wavelength shifts. As a benchmark, we consider the paper by Armani et al. [23], in which an ultrahigh quality factor ($Q>10^{8}$) whispering-gallery microcavity (WGM) was used for direct, label-free, single-molecule detection of interleukin-2 in serum. The WGM cavity was fabricated in a thermally oxidized SiO2 layer on top of a silicon wafer. The silica surface was functionalized to specifically bind the target molecule, with detection based on shifts in the resonant wavelength. Single-molecule sensitivity was confirmed by observing discrete binding events that induced measurable changes in the resonance frequency.
We now estimate the minimum detectable $\Delta n$ value that was observed in the paper and compare it with the expected values achievable using our proposed approach. In the benchmark paper, discrete resonance shifts on the order of $\sim$50 femtometers (fm) per single-molecule binding event were observed. The relationship between the resonance wavelength shift $\Delta\lambda$ and the change in the effective refractive index $\Delta n$ can be estimated as:
$\frac{\Delta\lambda}{\lambda_{r}}\approx\frac{\Delta n}{n_{\text{eff}}}.$ (18)
where $\lambda_{r}$ is the resonance wavelength (680 nm in the experiment) and $n_{\text{eff}}$ is the effective refractive index of the WGM mode. Rearranging Eq.  (18) to solve for $\Delta n$, we get $\Delta n\approx\left(\Delta\lambda/\lambda_{r}\right)\cdot n_{\text{eff}}$. Given that $\lambda_{r}=680$ nm and assuming $n_{\text{eff}}=2.5$, we estimate $\Delta n=\left(50\times 10^{-15}~{}\text{m}/680\times 10^{-9}~{}\text{m}\right)\cdot 2.5\approx 1.84\times 10^{-7}$. Therefore, the estimated change in the effective refractive index per single-molecule binding event is approximately $1\times 10^{-7}$ RIU. This is a rough estimation, but it gives the correct order of magnitude for $\Delta n$ achieved in the benchmark paper. This result has two important implications: 1) Achieving single-molecule detection is possible with a system that has a limit of detection (LOD) of at least $\sim 10^{-7}$ RIU; 2) Our proposed approach already achieves an LOD of $\sim 10^{-7}$ RIU with photonic cavities having $Q=10^{5}$ (see Table 2), which is at least three orders of magnitude lower in $Q$ compared to the $Q>10^{8}$ used in the benchmark study.
In addition, it is worth comparing the proposed approach with a commercially available alternative. Commercially available platforms such as Delta Life Science’s inQuiQ system [24] utilize microring resonator (MRR) structures to detect analyte-induced refractive index changes via resonance wavelength shifts, achieving state-of-the-art sensitivity with a baseline noise level of $\sim$0.01 RU (corresponding to $\sim 1\times 10^{-8}$ RIU) at a 1 Hz readout. According to publicly available information, this sensitivity is achieved using arrays of moderately high-$Q$ MRRs ($Q\approx 10^{5}$), operated near critical coupling to maximize the resonance shift per refractive index change $\Delta n$, and supported by active temperature stabilization and signal averaging to suppress baseline drift and noise. While highly effective, this approach relies on precise spectral tracking, which imposes integration challenges and demands high-resolution optical readout. It is unlikely that such systems employ on-chip integrated spectrometers, as achieving sub-picometer resolution with current CMOS-compatible technologies remains unfeasible; recent integrated spectrometer designs offer resolutions on the order of $\sim$100 pm [25], which is several orders of magnitude too coarse to resolve the required spectral shifts. In contrast, our platform employs a fundamentally different transduction mechanism based on TCSPC to monitor lifetime shifts of embedded quantum emitters in silicon, directly modulated by changes in the LDOS. Importantly, this lifetime-based approach enables comparable sensitivity with photonic cavities having the same $Q$ values ($10^{5}$$10^{6}$), significantly relaxing fabrication requirements. This makes our approach more scalable and robust, with inherent advantages in CMOS compatibility, noise resilience, and full on-chip integration without the need for spectral scanning or external tunable sources.
Given this, measuring the quantum emitter emission lifetime (via TCSPC) instead of tracking resonance wavelength shifts (via optical spectroscopy) offers several distinct practical and fundamental advantages:
  1. 1.
    Intrinsic sensitivity to LDOS changes: Lifetime directly reflects variations in the local density of optical states (LDOS). Even subtle environmental modifications can induce measurable lifetime shifts, potentially providing greater sensitivity than wavelength shifts — particularly for detecting very small refractive index changes.
  2. 2.
    Reduced complexity (no spectral measurements): Lifetime measurements are performed at a single wavelength and do not require wavelength-resolved spectroscopy. This enables simpler and more compact integration with on-chip photodetectors (e.g., SPADs), which are significantly cheaper and less complex than high-resolution optical spectrometers. Notably, achieving a resolution of 50 fm with on-chip integrated spectrometers remains unfeasible with current technology. For instance, one of the latest CMOS optoelectronic spectrometers based on photonic integrated circuits demonstrated a resolution of only $\sim$93 pm — approximately three orders of magnitude lower than what would be required for precise resonance tracking [25].
  3. 3.
    Lower susceptibility to temperature fluctuations: Emission lifetime is generally less affected by temperature-induced shifts compared to resonance wavelength, offering improved thermal stability and measurement reproducibility.
V Physical basis of the signal transduction amplifier
The above discussions indicate that this approach may be suitable for use without any effective signal transduction amplifier of the local refractive index change signal (such as an aptamer-functionalized hydrogel (aptagel) layer [26]), as it offers LODs on the order of $10^{-7}$$10^{-9}$ RIU, which is already potentially sufficient for single-molecule detection. However, it is nonetheless interesting to analyze the role of such a mechanical-optical amplifier and to visualize its potential impact on sensing performance. Eq. (13) can be rearranged as follows:
$\frac{\Delta n_{\text{opt}}}{\Delta\tau/\tau_{0}}=\frac{n_{\text{eff}}}{2Q}.$ (19)
Eq. (19) expresses that, for a given cavity (i.e., fixed $Q$ and $n_{\text{eff}}$), the relationship between refractive index changes $\Delta n$ and the measurable lifetime shift $\Delta\tau/\tau_{0}$ is fixed at the optimal operating point. Therefore, if one increases the actual refractive index shift $\Delta n$ — for example, due to aptagel collapse or shrinkage upon analyte binding — then the same sensing response (i.e., the same slope sensitivity) can be achieved with a higher $\Delta\tau/\tau_{0}$. In other words, while the aptagel does not alter the fundamental sensitivity of the photonic cavity (as governed by $Q$ and $n_{\text{eff}}$), it significantly increases the local refractive index change $\Delta n$ upon analyte binding. This, in turn, enables a proportional increase in the required measurable lifetime change $\Delta\tau/\tau_{0}$, thereby relaxing the timing resolution demands of the TCSPC detection system. In this way, the aptagel acts as an effective mechanical-optical amplifier of the analyte signal. To illustrate this concept, we can consider the following example: Without an aptagel, the binding of an analyte may induce a small refractive index change of approximately $1.2\times 10^{-7}$ RIU. For a photonic cavity with a quality factor $Q=10^{6}$, this corresponds to a relative lifetime change $\Delta\tau/\tau_{0}$ of about 0.3%. Detecting such a small change would require a time resolution of around 60 picoseconds (see Table 3) — a capability available in high-end TCSPC systems. In contrast, when a responsive aptagel is employed, the same analyte binding event could lead to a significantly larger refractive index change — for example, $1.2\times 10^{-6}$ RIU. This results in a relative lifetime change of approximately 3%, which would translate to a TCSPC absolute resolution requirement of $\sim$600 picoseconds. Such changes are much more easily detectable and can be resolved using standard, lower-cost TCSPC instrumentation. In this way, the aptagel does not increase the intrinsic sensitivity of the photonic cavity but instead relaxes the detection requirements by amplifying the local refractive index change resulting from analyte binding.
In addition, based on some realistic assumptions, we have attempted to “visualize” the effect of the aptagel in comparison to direct molecular binding to the sensor surface. When a single molecule binds directly to the surface of a SiO2 cavity, it perturbs the cavity’s optical mode through its local refractive index contribution. This perturbation is relatively small, as the molecule interacts only with the evanescent tail of the cavity mode. Let us assume the following: the refractive index of the molecule is $n_{\text{mol}}\approx 1.45$, the background medium has a refractive index of $n_{\text{bg}}\approx 1.33$, the volume of a single molecule is $V_{\text{mol}}\approx 10^{2}$ nm3, and the mode volume of the cavity is $V_{\text{mode}}\approx 10^{5}$ nm3. Then, the effective refractive index perturbation can be estimated as $\Delta n^{\text{direct}}=(n_{\text{mol}}-n_{\text{bg}})\cdot\left(V_{\text{mol}}/V_{\text{mode}}\right)\cdot f_{\text{field}}(\vec{r})$, where $f_{\text{field}}(\vec{r})$ is the normalized field intensity at the molecule’s position $\vec{r}$, relative to the mode field maximum. Assuming $f_{\text{field}}(\vec{r})=10^{-3}$, the resulting $\Delta n^{\text{direct}}\approx 1.2\times 10^{-7}$ RIU, which is in good agreement with our earlier estimation based on Ref. [23].
In the case of aptagel-based amplification, the effective refractive index perturbation can be estimated as:
$\Delta n^{\text{aptagel}}=\Delta\phi\cdot(n_{\text{gel}}-n_{\text{water}})\cdot f_{\text{overlap}},$ (20)
where $\Delta\phi$ is the increase in polymer volume fraction due to aptagel collapse (typically in the range 0.01–0.1), and $f_{\text{overlap}}$ is the field-weighted overlap between the collapsed gel and the cavity mode. The refractive indices of the gel and water are assumed to be $n_{\text{gel}}\approx 1.45$ and $n_{\text{water}}\approx 1.33$, respectively. Figure 4 shows the dependence of $\Delta n^{\text{aptagel}}$ on the polymer volume fraction increase $\Delta\phi$ for different values of $f_{\text{overlap}}$.
Figure 4: Dependence of $\Delta n^{\text{aptagel}}$ on the polymer volume fraction increase $\Delta\phi$ for different $f_{\text{overlap}}$ values.
At $\Delta\phi\rightarrow 0$, the curves closely approach the direct binding limit (dashed grey line), indicating that the aptagel has no amplification effect and behaves similarly to a single molecule bound to the surface. As $\Delta\phi$ increases, and for sufficiently high $f_{\text{overlap}}$ values ($10^{-3}$ or $10^{-4}$), even modest gel collapse results in a substantial increase in $\Delta n$, pushing the response well above the detection threshold and significantly facilitating measurement. This highlights the core advantage of aptagels — or any other amplification-capable functional layers: they can transform weak molecular binding signals into robust, volumetric refractive index changes.
VI Conclusion
We have presented a comprehensive theoretical framework for a new mechanism of probing refractive index changes based on Purcell-enhanced lifetime modulation of quantum emitters embedded in silicon photonic cavities. By operating the system off-resonance — at the point of maximum slope in the Purcell response — we identify a regime in which small refractive index perturbations produce large, linear, and readily measurable changes in fluorescence lifetime. This off-resonance strategy not only unlocks the full sensitivity potential of the system but also significantly relaxes the time-resolution requirements of the TCSPC instrumentation.
Our analysis demonstrates that, for quality factors in the range of $Q=10^{5}$$10^{7}$, the proposed sensing scheme achieves refractive index detection limits down to $10^{-9}$ RIU, with linear dynamic ranges suitable for single-molecule detection. Compared to conventional resonance-shift sensors, this lifetime-based approach offers several practical advantages: it is less susceptible to thermal and spectral noise, does not require high-resolution optical spectroscopy, and is fully compatible with integrated silicon photonic platforms. The use of long-lived emitters, such as T-centers, further enhances the system’s practicality by enabling high-$Q$ sensitivity to be accessed using standard, cost-effective TCSPC detectors.
Additionally, we show that incorporating functional layers such as aptagels can act as signal transduction amplifiers, further increasing the effective refractive index shift and thereby reducing the demands on TCSPC detection performance.
In summary, this work establishes a fundamentally different and technologically viable sensing modality — one that combines the robustness of quantum emitters with the scalability of silicon photonics and the practicality of lifetime-based detection. Although room-temperature operation of quantum emitters in silicon has yet to be demonstrated, the proposed platform offers a clear pathway toward next-generation, on-chip sensors with quantum-level sensitivity, operable at room temperature and without reliance on spectral resolution.
Acknowledgements.
No external funding was received for this research. The authors would like to thank the Armed Forces of Ukraine for providing security to perform this work.
Data Availability
The data that support the findings of this article are included within the article.